Mary Jo spends $1,860 to buy stock in two companies. She pays $15 a share to one of the companies and $26 a share to the other. If she ends up with a total of 80 shares, how many shares did she buy at $15 a share and how many did she buy at $26 a share?
She bought
shares at $15 and
shares at $26.
I am so confused!!!

Answer by fractalier(6550)

Okay, let me help you...
Let's let x be the number of shares she bought at $15.
Then the number of shares she bought at $26 must be 80-x.
The setup then looks like this:
15x + 26(80-x) = 1860
Now solve for x...we get
15x + 2080 - 26x = 1860
-11x = -220
x = 20 shares at $15 so that
80-x = 60 shares at $26 each
Now, that wasn't so bad, was it?